# Emotive Speech-to-Text Interfaces in XR: A Narrative Review of Psychophysiological and Accessibility Advances

### v2

@@ Abstract
This narrative review on emotional expression in Speech-to-Text (STT) interfaces with Virtual Reality (VR) aims to identify advancements, limitations, and research gaps in incorporating emotional expression into transcribed text generated by STT systems. Using a rigorous search strategy, relevant articles published between 2020 and 2024 are extracted and categorized into themes such as communication enhancement technologies, innovations in captioning, visual and affective augmentation, emotion recognition in AR and VR, and empathic machines. The findings reveal the evolution of tools and techniques to meet the needs of individuals with hearing impairments, showcasing innovations in live transcription, closed captioning, AR, VR, and emotion recognition technologies. Despite improvements in accessibility, the absence of emotional nuance in transcribed text remains a significant communication challenge. The study underscores the urgency for innovations in STT technology to capture emotional expressions. The research discusses integrating emotional expression into text through strategies like animated text captions, emojilization tools, and models associating emotions with animation properties. Extending these efforts into AR and VR environments opens new possibilities for immersive and emotionally resonant experiences, especially in educational contexts. The study also explores empathic applications in healthcare, education, and human-robot interactions, highlighting the potential for personalized and effective interactions. The multidisciplinary nature of the literature underscores the potential for collaborative and interdisciplinary research.
@@ 1 Introduction
This study embarks on a comprehensive exploration of the existing literature surrounding emotional expression in STT interfaces, with a specific emphasis on the psychophysiology aspect within Virtual Reality (VR). The objective is to identify advancements, limitations, and research gaps in the incorporation of emotional expression in transcribed text generated by STT systems. As communication technologies continue to advance, the ability to convey not only the semantic content but also the emotional nuances of speech becomes paramount for fostering richer and more meaningful interactions.
@@ 3.3 Emotion Recognition in AR and VR
More studies have discussed advancements in AR technologies for DHH individuals, particularly in the classroom setting [15]. The AR visual-captions system, implemented using Unity, ARKit, and AR Foundation, aims to provide real-time STT and visual elements around the teacher, creating an immersive learning experience. The study not only introduced the prototype but also outlined future research directions, especially on refining the design, conducting user studies, and extending the system’s capabilities for group conversations. Similarly, Real-time augmented reality visual-captions [16] for DHH children in classrooms builds on the AR-based system, providing a real-time solution for STT and keyword extraction, addressing the challenges faced by DHH children in mainstream education.
@@ 3.4 Speech Emotion Recognition: Techniques and Challenges
An important technique found in the literature is the use of color coding and visualization of emotion in speech, which offer a unique perspective on conveying emotional information. [25] utilizes haptic feedback vests and immersive virtual reality to quantify human emotion, providing insights for the development of emotionally relevant input devices. The use of bubble coloring in chat platforms [26, 27] explores the visual representation of speech emotion, considering color alterations and participant feedback for conveying specific emotions effectively. Other techniques in machine learning are being employed to include emotional recognition in text. One is Human-robot interaction and perception of emotions [28], demonstrating the effectiveness of animated text associated with emotional information in enhancing user interactions. The system’s potential applications in healthcare and education underscore its versatile impact.
Further, Schiano’s paper [29] delves into the perception of facial affect, bridging the gap between human facial expressions and prototype robot faces. The study informs the design of affective robot faces, contributing valuable insights into how emotions can be effectively communicated through facial expressions.In a related study,the analysis of humanoid avatar representations and emotions [30] delves into the uncanniness factor, highlighting the importance of selecting appropriate avatar types for accurate expression communication. The study’s findings support the need for avatar representations that effectively convey emotional cues in interactive systems. Finally, gestures can also be generated directly from speech using GANs [31] thereby opening up new possibilities in human-computer interaction. The user study using Turing test-inspired evaluation demonstrates the potential of the proposed technique in creating realistic gestures from speech. In conclusion, the narrative literature review traverses a spectrum of communication enhancement technologies, showcasing the evolution of tools and techniques to address the unique needs of deaf and hard of hearing individuals. From innovations in live transcription and closed captioning to advancements in augmented reality, emotive captioning, and speech emotion recognition, the literature reflects a dynamic landscape of research and development. The exploration of humanoid avatars, voice user interfaces, virtual reality, human-robot interaction, and the color coding of speech emotions further emphasizes the interdisciplinary nature of this field. The review also underscores the significance of understanding emotional expressions in text, images, and gestures, offering a holistic perspective on communication technologies that cater to diverse sensory and cognitive needs. As we navigate the intricacies of enhancing communication for individuals with hearing impairments, these technological advancements pave the way for a more inclusive and accessible future.
@@ 3.5 Empathic Machine
Empathic machine is also finding applications into wearable. In exploring the use of smart glasses equipped with an emotion recognition system to enhance human-to-human communication, with a focus on doctors and autistic adults [40], this study identifies the potential benefits for doctors in improving patient-doctor communication and for autistic adults facing challenges in adhering to neurotypical communication norms. The proposed system combines emotion recognition, AI, and smart glasses to provide real-time feedback on the emotional state of conversation partners. User evaluations indicate positive responses from both doctors and autistic adults, highlighting the potential for improving empathy and communication. Design considerations include customizable output preferences and addressing privacy and social impact concerns. The authors emphasize the system’s potential for educational purposes and ongoing development in the context of social impact and user experience. In a related study, [41] explores the use of virtual reality (VR) with wearable physiological sensors to examine how recalling autobiographical memories (AM) with emotional content affects an individual’s physiological state. By replicating the Autobiographical Memory Test (AMT) in VR, participants were presented with positive, negative, and neutral words to trigger memory recall. The study observed a positive influence of AM recall on electrodermal activity (EDA) peak amplitude, EDA peak number, and pupil diameter compared to situations without recall. However, emotional AM recall did not produce a significant impact. The article concludes by discussing the potential and limitations of utilizing autobiographical memory to enhance personalized mobile VR experiences in conjunction with physiological sensors. The findings reveal a significant effect of AM recall on EDA mean peak amplitude, EDA peak number, and pupil diameter compared to a no-recall condition. However, no significant differences were identified in emotional AM recall (positive, negative, neutral).
Visual and Affective Augmentation in Captioning
@@ 4 Discussion
This review explores advancements, limitations, and gaps in the incorporation of emotional expression in STT interfaces, particularly within immersive environments such as VR and AR. While STT technologies have significantly improved accessibility—especially for DHH users—most systems lack affective subtleties, reducing the depth and naturalness of communication. The review was motivated by the critical absence of emotional expression in traditional captioning systems.
The review was conducted to address a critical aspect overlooked in traditional language transcription services—emotional expression. While STT systems enhance accessibility for the DHH, the loss of emotional nuance in transcribed text poses a communication challenge. The review seeks to analyze and synthesize available literature to understand the current state of research, advancements, and gaps in reintegrating emotional expression into transcribed text.
Through an extensive literature review spanning databases such as ACM Digital Library, IEEE Explore, Scopus, and Google Scholar, the study identified 31 relevant papers. The review explored communication enhancement technologies, innovations in captioning, emotion recognition in AR and VR, and empathic machines. The review successfully organized the literature into thematic categories, identified trends, and presented key findings. It critically discussed the implications of the literature on the research objective. Notably, a thematic synthesis on visual and affective augmentation was added to highlight recent progress in emotion visualization within XR environments. These studies show promise in supporting emotional expression through ambient cues and interactive visual metaphors. However, they remain disconnected from real-time STT outputs, underscoring the continued lack of emotionally expressive captions. Figure 1 provides a high-level overview of the thematic areas synthesized in this review, showing how emerging approaches in XR-based captioning intersect with accessibility, psychophysiology, and affective computing.
The limited relevant studies found highlights the need for more research in this area. Also, of the relevant papers found, there is no study that addressed incorporating emotional expression in STT in an AR or VR environment, thus further supporting more research in this area.
The main findings of the narrative review highlight the evolution of tools and techniques aimed at addressing the unique needs of individuals with hearing impairments and other users of language transcription text. This includes recent advances in visual and affective augmentation strategies that address the cognitive and emotional challenges of caption use in immersive XR environments.
Notably, emerging research in affective visualization—such as emotion-driven caption overlays, color-coded text, and bi-directional feedback systems—represents a critical shift toward personalized and empathetic captioning. These approaches, which blend psychophysiology with design, offer promising pathways to reduce cognitive load and increase emotional resonance for both DHH and general users in XR spaces.
However, the notable absence of emotional expression in STT captioning applications remains a significant limitation, as highlighted by [9]. Despite their commendable role in enhancing accessibility, conventional captioning systems often fail to convey affective subtleties such as tone, emphasis, or emotional context [6], limiting the richness and expressiveness of communication [5]. While recent studies have explored visual representations of emotion in XR environments [20, 23, 19], these efforts are primarily focused on ambient or avatar-mediated emotional cues rather than direct integration into ASR-generated captions. This reveals a critical research gap: the lack of emotionally expressive, real-time captions in immersive environments. Bridging this gap requires further investigation into how emotional visualization techniques can be tightly coupled with STT outputs to improve both comprehension and engagement for DHH users and others relying on captions in XR.
The exploration of strategies to integrate emotional expression into text, such as Rashid’s framework [7, 5], signifies a shift beyond traditional captioning methods. Animated text captions, enriched with graphics and color, not only show positive responses among hard of hearing viewers [8] but also present a potential avenue for more inclusive and emotionally expressive communication.
The extension of these efforts into Augmented Reality (AR) and Virtual Reality (VR) environments, where technologies offer unique advantages [12], opens new possibilities for immersive and emotionally resonant experiences. While studies dedicated to enhancing STT captions in AR and VR environments are promising [16, 15], the results underscores the need for continued research to infuse emotional expression into these advanced platforms, particularly in educational training contexts [18, 17].
The implementation of the emojization tool [9], despite its limitations in covering a sufficient vocabulary, demonstrates a tangible effort to restore emotion to text in the STT transcription process. This example serves as a valuable step toward addressing the emotional gap in communication technologies, particularly within the AR/VR environment.
Furthermore, the application of emotional recognition through machine learning and robotics in healthcare and education [27, 28, 29, 30] highlights the broader impact of these results. The effective communication of emotions in interactive systems has implications for personalized and empathetic interactions, potentially improving patient care and educational experiences.
The digitization of human emotions, as demonstrated in studies such as Roshdy et al. [37], has direct implications for healthcare applications. The development of Brain-Computer Interface (BCI) systems capable of interpreting and responding to human emotions opens new possibilities for enhancing patient care and interaction with humanoid robots, suggesting a potential revolution in healthcare practices.
The exploration of models like the Automatic Cognitive Empathy Model (ACEM) for humanoid robots [39] contributes to more extended and engaging human-robot interactions. This has implications for various fields, including customer service, education, and entertainment, where robots can adapt their behavior based on users’ emotional states, leading to more personalized and effective interactions.
Wearable technologies, such as smart glasses with emotion recognition systems [40], have the potential to significantly improve communication in specific contexts, such as doctor-patient interactions and communication challenges faced by autistic adults. The real-time feedback on emotional states can lead to more empathetic and tailored communication, addressing specific needs in these domains.
The discussed systems, particularly those involving wearable technologies and intelligent interfaces, have implications for educational settings. The potential for improving empathy and communication, as well as considerations for customizable output preferences, privacy, and social impact, suggests that these technologies could play a role in educational environments, fostering better communication and understanding.
The exploration of virtual reality (VR) experiences with wearable physiological sensors [41] provides insights into the potential and limitations of leveraging autobiographical memory to enhance personalized mobile VR experiences. This understanding is valuable for developers and researchers aiming to create immersive and emotionally resonant virtual experiences.
The multidisciplinary nature of these studies, incorporating fields such as neuroscience, artificial intelligence, and psychology, highlights the potential for collaborative and interdisciplinary research. This can lead to the development of more holistic and effective empathic systems that consider both technological and human factors.
The significance of this research lies in its contribution to advancing communication technologies for individuals with hearing impairments. By exploring emotional expression in STT interfaces, the study sheds light on the evolving landscape of tools and techniques. The research emphasizes the importance of inclusivity and accessibility in communication technologies, fostering a more empathetic and inclusive future.
@@ 4.2 Gaps, Challenges and Opportunities
    Innovations in emotive captioning and speech emotion recognition warrant further exploration, particularly in educational setting where DHH students, including other categories of students rely on STT as their main method of communication during lectures.
@@
    Explore the integration of cognitive empathy and improve emotion detection in short conversations. The study’s promising results underscore the potential for empathy-driven enhancements in medical chatbots [33].
@@
    Prioritize the development of an automatic metric for assessing empathy in text-based messages, eliminating reliance on human judges and enhancing result comparability. Refining the definition of empathy to encompass additional aspects like nationality and educational level is another crucial step. 2. Further research should incorporate finer emotion categories and measure emotion intensity. Improving the dialogue system’s adaptability to learn, unlearn, and re-learn from user interactions, monitoring emotional states and personal information, is essential. 3. Additionally, addressing ethical concerns related to enhanced empathy and human-like qualities in chatbots, including guidelines on handling user information and drawing conclusions, should be a focus of future investigations [35].
@@
    Explore alternative markers for empathic responses, involve different character styles and modes of realism, and incorporate physiological or behavioral measures of empathy. 2. Additionally, replicating the study in a virtual reality context was suggested to further understand the role of immersive environments in influencing empathic experiences with virtual characters [42].
@@
    Future work to incorporate video and audio data with EEG and physiological signals for a comprehensive approach to emotion recognition in conversations. The methods and dataset presented are anticipated to benefit various intelligent user interface applications [38].
@@
    Explore the system’s potential applications in educational settings, particularly for training and educating medical professionals. Assess how the system could be integrated into medical education curricula [40].
@@
    Future research with a larger participant pool, the inclusion of higher emotional intensity words, and exploration of additional physiological signals like EEG and heart rate variability (HRV). The study’s implications extend to designing text-based VR interfaces, passive learning, exposure treatment, and personalized VR gaming experiences [41].
@@
    Explore the integration of real-time emotional visualization techniques with ASR captions in XR. While recent studies demonstrate emotion-aware avatars, ambient indicators, and user-controlled visual augmentations, these are not yet linked to STT outputs, revealing a critical opportunity to design emotionally expressive captioning systems.
4.3 Next Steps
This study has revealed that research in this domain is greatly limited, and this is evident in the few papers deemed eligible for inclusion (only 31 papers). Based on the available literature reviewed, there is limited study on the subject of incorporating emotional expression in STT in AR/VR for use in educational setting purposes. Therefore, it is necessary to conduct further research to fill the identified gap. To do this, my next step will involve working with the following research questions as my Ph.D. research work:

### v3

# Narrative Review of Emotional Expression Support in XR: Psychophysiology of Speech-to-Text Interfaces

@@ Abstract
This narrative review examines recent advancements, limitations, and research gaps in integrating emotional expression into speech-to-text (STT) interfaces within extended reality (XR) environments. Drawing from 37 peer-reviewed studies published between 2020 and 2024, we synthesized literature across multiple domains—including affective computing, psychophysiology, captioning innovation, and immersive human-computer interaction. Thematic categories include communication enhancement technologies for Deaf and Hard of Hearing (DHH) users, emotive captioning strategies, visual and affective augmentation in AR/VR, speech emotion recognition, and the development of empathic systems. Despite the growing accessibility of real-time STT tools, such systems largely fail to convey affective nuance, limiting the richness of communication for DHH users and other caption consumers. This review highlights emerging approaches such as animated captions, emojilization, color-coded overlays, and avatar-based emotion visualization, but finds a persistent gap in real-time emotion-aware captioning within immersive XR contexts. We identify key research opportunities at the intersection of accessibility, XR, and emotional expression, and propose future directions for the development of affect-responsive, user-centered captioning interfaces.
@@ 1 Introduction
This study embarks on a comprehensive exploration of the existing literature surrounding emotional expression in STT interfaces, with a specific emphasis on the psychophysiology aspect within Extended Reality (XR). The objective is to identify advancements, limitations, and research gaps in the incorporation of emotional expression in transcribed text generated by STT systems. As communication technologies continue to advance, the ability to convey not only the semantic content but also the emotional nuances of speech becomes paramount for fostering richer and more meaningful interactions.
@@ 3.3 Emotion Recognition in AR and VR
More studies have discussed advancements in AR technologies for DHH individuals, particularly in the classroom setting [15]. The AR visual-captions system, implemented using Unity, ARKit, and AR Foundation, aims to provide real-time STT and visual elements around the teacher, creating an immersive learning experience. The study not only introduced the prototype but also outlined future research directions, especially on refining the design, conducting user studies, and extending the system’s capabilities for group conversations. Similarly, Real-time AR visual-captions [16] for DHH children in classrooms builds on the AR-based system, providing a real-time solution for STT and keyword extraction, addressing the challenges faced by DHH children in mainstream education.
@@ 3.4 Speech Emotion Recognition: Techniques and Challenges
An important technique found in the literature is the use of color coding and visualization of emotion in speech, which offer a unique perspective on conveying emotional information. [25] utilizes haptic feedback vests and immersive VR to quantify human emotion, providing insights for the development of emotionally relevant input devices. The use of bubble coloring in chat platforms [26, 27] explores the visual representation of speech emotion, considering color alterations and participant feedback for conveying specific emotions effectively. Other techniques in machine learning are being employed to include emotional recognition in text. One is Human-robot interaction and perception of emotions [28], demonstrating the effectiveness of animated text associated with emotional information in enhancing user interactions. The system’s potential applications in healthcare and education underscore its versatile impact.
Further, Schiano’s paper [29] delves into the perception of facial affect, bridging the gap between human facial expressions and prototype robot faces. The study informs the design of affective robot faces, contributing valuable insights into how emotions can be effectively communicated through facial expressions.In a related study,the analysis of humanoid avatar representations and emotions [30] delves into the uncanniness factor, highlighting the importance of selecting appropriate avatar types for accurate expression communication. The study’s findings support the need for avatar representations that effectively convey emotional cues in interactive systems. Finally, gestures can also be generated directly from speech using GANs [31] thereby opening up new possibilities in human-computer interaction. The user study using Turing test-inspired evaluation demonstrates the potential of the proposed technique in creating realistic gestures from speech. In conclusion, the narrative literature review traverses a spectrum of communication enhancement technologies, showcasing the evolution of tools and techniques to address the unique needs of DHH individuals. From innovations in live transcription and closed captioning to advancements in AR, emotive captioning, and speech emotion recognition, the literature reflects a dynamic landscape of research and development. The exploration of humanoid avatars, voice user interfaces, VR, human-robot interaction, and the color coding of speech emotions further emphasizes the interdisciplinary nature of this field. The review also underscores the significance of understanding emotional expressions in text, images, and gestures, offering a holistic perspective on communication technologies that cater to diverse sensory and cognitive needs. As we navigate the intricacies of enhancing communication for individuals with hearing impairments, these technological advancements pave the way for a more inclusive and accessible future.
@@ 3.5 Empathic Machine
Empathic machine is also finding applications into wearable. In exploring the use of smart glasses equipped with an emotion recognition system to enhance human-to-human communication, with a focus on doctors and autistic adults [40], this study identifies the potential benefits for doctors in improving patient-doctor communication and for autistic adults facing challenges in adhering to neurotypical communication norms. The proposed system combines emotion recognition, AI, and smart glasses to provide real-time feedback on the emotional state of conversation partners. User evaluations indicate positive responses from both doctors and autistic adults, highlighting the potential for improving empathy and communication. Design considerations include customizable output preferences and addressing privacy and social impact concerns. The authors emphasize the system’s potential for educational purposes and ongoing development in the context of social impact and user experience. In a related study, [41] explores the use of VR with wearable physiological sensors to examine how recalling autobiographical memories (AM) with emotional content affects an individual’s physiological state. By replicating the Autobiographical Memory Test (AMT) in VR, participants were presented with positive, negative, and neutral words to trigger memory recall. The study observed a positive influence of AM recall on electrodermal activity (EDA) peak amplitude, EDA peak number, and pupil diameter compared to situations without recall. However, emotional AM recall did not produce a significant impact. The article concludes by discussing the potential and limitations of utilizing autobiographical memory to enhance personalized mobile VR experiences in conjunction with physiological sensors. The findings reveal a significant effect of AM recall on EDA mean peak amplitude, EDA peak number, and pupil diameter compared to a no-recall condition. However, no significant differences were identified in emotional AM recall (positive, negative, neutral).
@@ 4 Discussion
This review identifies key advancements, limitations, and research gaps in the integration of emotional expression into speech-to-text (STT) interfaces, particularly within immersive environments such as virtual and augmented reality (VR/AR). While STT technologies have significantly improved accessibility—especially for DHH users—they typically fail to convey affective cues, resulting in reduced expressiveness and diminished communication quality [6]. Motivated by this persistent limitation in traditional captioning systems, our review synthesizes current literature to assess the state of research and inform future directions for more emotionally resonant captioning systems.
Through a structured narrative review of 37 peer-reviewed studies published between 2020 and 2024, we examined developments across communication enhancement technologies, captioning innovation, emotion recognition in XR, and empathic human-computer systems. Our thematic synthesis included a focused review of visual and affective augmentation strategies, which show promise for conveying emotion through ambient cues and interactive design elements [20, 19, 23]. However, these strategies are largely decoupled from real-time STT outputs, underscoring a significant gap in current systems: the absence of emotionally expressive, real-time captions in immersive environments.
Despite growing interest in affective computing and XR accessibility, few studies directly address the integration of emotion into STT interfaces within AR/VR contexts. Current approaches focus primarily on visual enhancements or emotion-aware avatars, without embedding affective cues directly into the transcribed text [9, 5]. This limitation presents a critical opportunity for innovation—particularly in education and training scenarios where emotional tone can support comprehension, reduce cognitive load, and increase learner engagement [17, 18].
Research into captioning innovations such as Rashid’s animated text framework [7, 5] and Hu’s emojization tool [9] exemplifies early efforts to embed emotion into transcription. These methods—whether through animated captions, color cues, or emoji augmentations—have shown positive responses from DHH users [8], and highlight the potential for more expressive, inclusive STT systems. Yet, scalability and vocabulary limitations remain challenges, especially for deployment in real-time XR applications.
Extending these innovations into immersive platforms offers new possibilities. Recent AR and VR captioning prototypes demonstrate technical feasibility for real-time STT with spatial awareness and keyword extraction [16, 15]. However, few efforts have integrated emotional nuance into these systems. In educational settings, where XR-based captioning is increasingly applied, the inclusion of affective information could enhance both accessibility and engagement [17, 18]. Research on presence and affective response in VR environments further suggests that immersive platforms are well-suited for emotionally resonant experiences [12].
Beyond captioning, affective computing in adjacent domains—such as healthcare, robotics, and human-agent interaction—provides valuable insights. Studies show that emotionally responsive systems, powered by machine learning, facial recognition, and physiological sensing, can enhance empathy, personalization, and user satisfaction [28, 27, 30]. Applications in education and clinical settings demonstrate that recognizing and responding to user emotion improves outcomes and fosters trust [29]. The digitization of human emotion, including efforts using brain-computer interfaces (BCIs) and affective robotics, represents a broader trend toward emotionally intelligent technology. BCI-driven systems that interpret EEG or physiological signals offer promising methods for capturing emotional states [37, 38], though their integration with STT remains underexplored. Similarly, empathy models like ACEM for humanoid robots illustrate how emotional awareness can be incorporated into interaction design to enable adaptive and human-like behavior [39].
Wearable technologies, such as smart glasses with built-in emotion recognition, have also shown promise in improving communication among autistic individuals and between doctors and patients [40]. These tools deliver real-time emotional feedback to enhance interpersonal understanding, and their integration into educational contexts could support more personalized and empathetic learning experiences. Multisensory XR systems that combine physiological data, autobiographical memory, and real-time feedback offer new frontiers for emotionally responsive interaction. Studies exploring such experiences emphasize the potential for tailoring virtual environments to individual affective states [41], providing valuable frameworks for the design of emotionally expressive captioning systems.
Overall, the findings of this review suggest a compelling case for interdisciplinary research that bridges affective computing, psychophysiology, XR design, and accessibility. The future of inclusive STT systems lies in integrating emotional expression not as a secondary visual layer, but as an intrinsic component of the transcription process—transforming captions from neutral information displays into emotionally meaningful communication tools.
In sum, advancing emotional expression in STT systems holds particular promise for enhancing communication accessibility for DHH users. By embedding emotion-aware cues into real-time captions, especially within XR environments, designers and researchers can move toward more human-centered, empathetic technologies. This work underscores the importance of inclusivity not only in functionality but in emotional fidelity—ensuring that all users, regardless of hearing ability, can access the full depth of spoken communication.
@@ 4.2 Gaps, Challenges and Opportunities
    Innovations in emotive captioning and speech emotion recognition warrant further exploration, particularly in educational settings where DHH students and other learners rely on STT as their main method of communication during lectures.
@@
    Further integration of cognitive empathy models is needed to improve emotion detection in short conversations. Enhancements in medical chatbots show potential for empathy-driven systems [33].
@@
    There is a need for automated metrics to assess empathy in text-based systems. Future studies should also refine definitions of empathy and explore emotion intensity and categories, while addressing ethical concerns in human-like chatbot design [35].
@@
    Research should explore alternative empathic response markers, including diverse character styles and physiological indicators, and test these in immersive environments like VR [42].
@@
    Emotion recognition systems should incorporate multimodal inputs, including EEG and physiological signals, for a holistic approach to understanding affective states in conversation [38].
@@
    Future investigations should assess applications of emotion-aware systems in educational contexts, including medical education, and evaluate their integration into training curricula [40].
@@
    Larger-scale studies are needed with more emotionally intense stimuli and diverse biosignals, applicable to VR learning, exposure therapy, and adaptive gaming [41].
@@
    The field should prioritize real-time integration of emotional visualization techniques with ASR captions in XR, moving beyond ambient indicators to emotion-aware captions that directly reflect STT output.
4.3 Future Work Directions
To address the outlined challenges and research gaps, we propose the following directions for future exploration: